OpenStudy (anonymous):

HELP!! Will FAN AND MEDAL! ALGEBRA 1!! PLEASE!! The Jackson High School gymnastics team sells calendars for their annual fundraiser. The function rule below describes the amount of money the team can raise, where y is the total amount in dollars, and x is the number of calendars sold. y=3x+28 Use the function rule to find the corresponding range values when the indicated number of calendars is sold. Show your work for finding each range value. a. 50 calendars b. 75 calendars c. 90 calendars d. Create a table of values that includes the x- and y-values for parts a-c. Write a sentence to explain what the range represents for this func

OpenStudy (anonymous):

So I've figured out the first portion but I am still confused. I need a bit of help with the next portion please. Here's what I have so far... In the function rule, I replace x with 50 in y = 3x + 28 because 3x means 3 times x so I multiplied 50 by 3, when I received that answer, I add it by 28. That will give me the amount of money the gymnastic team will raise by selling 50 calendars. y = 3x + 28 y = (3 x 50)(150 + 28) y = 150 + 28 = $178 y = $178 total